OpenStudy (anonymous):

How many solutions does the equation −4y + 4y + 2 = 2 have? Two None One Infinitely many

OpenStudy (anonymous):

@Janu16 @KendrickLamar2014 @kropot72

OpenStudy (anonymous):

@Mehek14 @Nurali @texaschic101

OpenStudy (anonymous):

oh sorry I was thinking none

OpenStudy (anonymous):

hey I guess I solved it wrong sorry

OpenStudy (anonymous):

think*

OpenStudy (anonymous):

all real numbers... - courtesy of google and aloneS

alones (alones):

Okay so −4y+4y+2=2 −4y+4y+2=2 simplify (−4y+4y)+(2)=2(Combine Like Terms) 2=2 don't forget to ubtract 2 from both sides. 2−2=2−2 0=0 I believe that all the real numbers are solutions.

alones (alones):

AloneS? what about it?

OpenStudy (anonymous):

thanks great ;D I have a few more questions so I'll re-post lol